A particular fruit's weights are normally distributed, with a mean of 784 grams and a standard deviation of 10 grams. The heaviest 17% of fruits weigh more than how many grams?

Give your answer to the nearest gram.
Answer #1

Solution :

Given that,  

mean = \mu = 784

standard deviation = \sigma = 10

Using standard normal table ,

P(Z > z) = 17%

1 - P(Z < z) = 0.17

P(Z < z) = 1 - 0.17 = 0.83

P(Z < 0.95) = 0.83

z = 0.95

Using z-score formula,

x = z * \sigma + \mu

x = 0.95 * 10 + 784 = 794

The heaviest 17% of fruits weigh more than is 794 grams